OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. 1. (1 pt) What property is shown by the equation? Column A Column B 1. (34 + 11) + 9 = (11 + 34) + 9 2. 25 + (–15 + 4) = (25 + (–15)) + 4 3. –(–21 + 8) = 21 + (–8) 4. –30 + (–16) = –16 + (–30) A. commutative property B. associative property C. opposite of a sum property

OpenStudy (anonymous):

1234 is column a and abc is column b

OpenStudy (anonymous):

okay do you know what the properties look like?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay so the commutative property is: a + b = b + a

OpenStudy (anonymous):

associative property: a+(b+c)=(a+b)+c

OpenStudy (anonymous):

thats 2 properties shown, you see any matches?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

help im bad at math

OpenStudy (anonymous):

its all good man.... dont worry.. well lets do commutative first. so a+b=b+a... -30+(-16)=-16+(-30) i did the first one... now you try and find the associative property..

OpenStudy (anonymous):

so if the associative property is: a+(b+c)=(a+b)+c

OpenStudy (anonymous):

which equation is similar?

OpenStudy (anonymous):

@ArrmaTyphon